how can u solve f(x)=2x2squared-4
Answer by gonzo(654)

if i understand you correctly, you want to solve f(x) = 2*x^2 - 4 when x = 2.
this becomes f(2) = 2*2^2 - 4
this becomes f(2) = 4 - 4
this becomes f(2) = 0
answer would be f(2) = 0
if i understand you correctly.